hey,
could answer this question for me.
find the quadratic relation in vertex form that has zeros -3 and 5 and passes through (3,6).
i found the x-symmetry which is -3+5= 2/2 = 1
Therefore x=1
but i did get how to find the y value
could u please solve it to me and send it back.
i will be waiting.
thank you,

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

find the quadratic relation in vertex form that has zeros -3 and 5 and passes through (3,6).
------------------
f(x) = a(x+3)(x-5)
f(3)=6
Therefore:
f(3) = a(3+3)(3-5)=6
a*6*-2=6
-12a = 6
a = -1/2
----------------
EQUATION:
f(x) = (-1/2)(x+3)(x-5)
---------------
This quadratic has zeros at x=-3,and x=5, and passes thru (3,6)
============================================================
Cheers,
Stan H.

Answer by edjones(8007)

I goofed.
Stanbon did it the correct way.
I hope I didn't confuse you any.
Ed